# Disentangled Representation Learning for Parametric Partial Differential Equations

### v1

@@ B.2 Supervised forward and inverse PDE learning
  • MetaNO: We use a 16-layer IFNO model with the lifting layer as the adaptive layer. We keep the total number of parameters in MetaNO the same as the number of parameters used in forward PDE learning in DisentangO.
@@
  • NIO: We closely follow the setup in Molinaro et al. (2023) and use four convolution blocks as the encoder for the branch net and a fully connected neural network with two hidden layers of 256 neurons as the trunk net, with the number of basis functions set to 50. For the FNO part, we use one Fourier layer with width 32 and modes 8, as suggested in Molinaro et al. (2023).
@@
  • FNO: We adopt a 4-layer FNO with width 26 and modes 8. For the forward model, in addition to the loading field and coordinates as input, we also concatenate the ground-truth material properties to form the final input. For the inverse model, we employ an iterative gradient-based optimization to solve for the optimal material parameters for physics discovery.
@@ B.3 Mechanical MNIST benchmark
Herein, we demonstrate the latent traversal based on a randomly picked loading field as DisentangO-2 input. One can clearly see that the digit changes from “6” to “0” and then “2” from the top left moving down, and from “6” to “1” and then “7” moving to the right. Other digits are visible as well such as “7”, “9”, “4” and “8” in the right-most column. This corresponds well to the distribution of the latent clustering in Figure 3. We also provide two exemplary MMNIST latent interpretations of the learned DisentangO-15 in Figures 7 and 8.
@@
Figure 7: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 10 encodes the information for digit ‘8’, as is evidenced by the fact that all $y<-0.4$ regions on the scatterplots contain only digit ‘8’.
@@
Figure 8: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 12 encodes the information for digit ‘3’, as is evidenced by the fact that all $y<-0.45$ regions on the scatterplots contain only digit ‘3’.
@@ Appendix C Data Generation
C.1 Example 1 - Supervised forward and inverse PDE learning
@@
$+\frac{k_{1}}{2k_{2}}(e^{k_{2}\langle S(\alpha)\rangle^{2}}+e^{k_{2}\langle S(-\alpha)\rangle^{2}}-2)+\frac{E}{6(1-2\nu)}(\frac{J^{2}-1}{2}-\ln J)\text{ ,}$ (C.1)
@@
C.2 Example 2 - Mechanical MNIST benchmark
@@
C.3 Example 3 - Unsupervised heterogeneous material learning
@@
Varying fiber orientation distribution. We generate the samples by controlling the predefined parameters of Gaussian Random Field (GRF). With the GRF sharpened by the thresholding, the values to binary field represent two distinct fiber orientations. The binary field is smoothed using a windowed convolution. To address boundary conditions, the matrix is padded with replicated edge values, ensuring that the convolution works uniformly across the entire grid. After the smoothing process, the padded sections are removed, and the remaining field is used to construct the fiber field. We use two fixed fiber angles $\frac{\pi}{3},\frac{2\pi}{3}$ for the corresponding binary field. We generate 300 material sets, each with 500 loading/displacement pairs, and divide these into training, validation, and test sets in a 200/50/50 split. Exemplar Samples from this dataset are illustrated in Figure 11. These samples demonstrate the variability of $\bm{b}$, which is of critical for the latent variable identifiability in unsupervised learning settings, as proved in Theorem 2.
Varying fiber orientation magnitude and segmentation line rotation. Instead of controlling the fiber orientation angles on the binary field as two constant values, we sample the orientation distribution consisting of two segments with orientations $\alpha_{1}$ and $\alpha_{2}$ on each side, respectively, separated by a line passing through the center. The values of $\alpha_{1}$ and $\alpha_{2}$ are independently sampled from a uniform distribution over $[0,2\pi]$, and the centerline’s rotation is sampled from $[0,\pi]$. We generate 300 material sets, each with 500 loading/displacement pairs, and divide these into training, validation, and test sets in a 200/50/50 split.
@@
The loading in this example is taken as the body load, $\bm{f}(\bm{x})$. Each instance is generated as the restriction of a 2D random field, $\phi(\bm{x})=\mathcal{F}^{-1}(\gamma^{1/2}\mathcal{F}(\Gamma))(\bm{x})$. Here, $\Gamma(\bm{x})$ is a Gaussian white noise random field on $\mathbb{R}^{2}$, $\gamma=(w_{1}^{2}+w^{2}_{2})^{-\frac{5}{4}}$ represents a correlation function, $w_{1}$ and $w_{2}$ are the wave numbers on $x$ and $y$ directions, respectively, and $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and its inverse, respectively. This random field is anticipated to have a zero mean and covariance operator $C=(-\Delta)^{-2.5}$, with $\Delta$ being the Laplacian with periodic boundary condition on $[0,2]^{2}$, and we then further restrict it to $\Omega$. For the detailed implementation of Gaussian random field sample generation, we refer interested readers to Lang and Potthoff (2011a). Then, for each sampled loading $\bm{f}_{i}(\bm{x})$ and microstructure field $\bm{b}^{\eta}(\bm{x})$, we solve for the displacement field $\bm{u}^{\eta}_{i}(\bm{x})$ on the entire domain.

### v2

@@ B.2 Supervised forward and inverse PDE learning
  • MetaNO: We use a 8-layer IFNO model with the lifting layer as the adaptive layer. We keep the total number of parameters in MetaNO the same as the number of parameters used in forward PDE learning in DisentangO.
@@
  • NIO: We closely follow the setup in Molinaro et al. (2023), where two neural operators (DeepONet and FNO) are stacked together to realize the operator-to-function intuition. The first operator maps multiple solution functions to a set of representations (which can be seen as an analog of eigenfunctions), and the second operator infers the underlying parameter field from the mixed representations. As NIO requires the solution field as input, it cannot be used as a forward solver. Hence, NIO only solves the inverse PDE problem, and it can only be applied to the fully supervised setting. Specifically, we use four convolution blocks as the encoder for the branch net and a fully connected neural network with two hidden layers of 256 neurons as the trunk net, with the number of basis functions set to 50. For the FNO part, we use one Fourier layer with width 32 and modes 8, as suggested in Molinaro et al. (2023).
  • FNO: Since FNO is originally designed as a function-to-function mapping, we consider the inverse optimization procedure following Lee et al. (2024), and develop a two-phase process to solve the forward and inverse problems sequentially. In the first phase, we construct the forward mapping from the loading field $\bm{f}$ and the ground-truth material parameter $\bm{b}$ to the corresponding solution $\bm{u}$ as: $\mathcal{G}^{FNO}[\bm{f},\bm{b};\theta^{FNO}](\bm{x})=\bm{u}(\bm{x})$. This can be seen as an analog of the forward solution operator $\mathcal{G}$ in our setting. Then, with the trained FNO as a surrogate for the forward solution operator, we fix its NN parameters $\theta^{FNO}$, and use it together with gradient-based optimization to solve for the optimal material parameters as an inverse solver. Specifically, given a set of loading/solution data pairs $\{(f_{i},u_{i})\}_{i=1}^{N}$, we start from a random guess of the underlying material parameters (typically chosen as the average of all available instances of material parameters for fast convergence), and minimize the difference between the predicted displacement field from FNO and the ground-truth one:
    $b^{*}=\text{argmin}_{b}\sum_{i=1}^{N}\|u_{i}-\mathcal{G}^{FNO}[f_{i},b;\theta^{FNO}]\|^{2}.$
@@
    As the FNO parameters are fixed, we can back propagate this loss and optimize the input material parameters in an iterative fashion. We adopt a 4-layer FNO with width 26 and modes 8. For the forward model, in addition to the loading field and the coordinates as input, we also concatenate the ground-truth material properties to form the final input. For the inverse model, we employ an iterative gradient-based optimization to solve for the optimal material parameters for physics discovery. For fair comparison, in terms of the averaged per-epoch runtime, we report the sum of both the forward and inverse solvers. The averaged per-epoch runtimes for the forward solver and the inverse solver are 6.2 seconds and 2.9 seconds, respectively, accounting for the total per-epoch runtime of 9.1 seconds in Table 1.
  • UFNO: The 2D U-FNO model extends the Fourier Neural Operator (FNO) architecture by incorporating U-Net-style skip connections. The network consists of six spectral convolution blocks, each combining a global Fourier operator and a local $1\times 1$ convolution. In the later three layers, additional skip-enhanced feature extraction is provided by U-Net blocks. The input consists of spatial features $f(x,y)$, concatenated with coordinate embeddings $(x,y)$ and 5 material parameters. The input is lifted to a higher-dimensional latent space via a linear layer of size $[8,20]$. Each of the six layers performs the following composite operation:
    $v_{j+1}=\text{ReLU}(\mathcal{F}_{j}(v_{j})+\mathcal{W}_{j}(v_{j})+\mathcal{U}_{j}(v_{j})),$
    where $\mathcal{F}_{j}$ is a spectral convolution layer (2D Fourier transform, mode truncation, linear transformation, and inverse transform), $\mathcal{W}_{j}$ is a $1\times 1$ convolution, and $\mathcal{U}_{j}$ is a U-Net block (included only in layers 4 to 6). After all six layers, the output is projected through two fully connected layers of sizes $[20,128]$ and $[128,2]$ to produce the final 2-channel output.
  • WNO: The Wavelet Neural Operator (WNO) consists of 4 stacked wavelet kernel integral layers. The input is first lifted to a high-dimensional representation using a fully connected layer of size $[8,27]$. Each of the 4 layers combines a learned local operator $W$ (implemented as $1\times 1$ convolution) with a non-local wavelet-based integral operator $K$ (via continuous wavelet transform). The architecture follows: The input $f(x,y)$ is first augmented by concatenating it with the positional grid coordinates $(x,y)$ and additional 5 material parameters. This augmented input is then lifted into a higher-dimensional space via a fully connected layer that maps the input to a hidden width of width. The core of the architecture consists of 4 wavelet-based layers, the output is projected back to the desired output space using two linear layers of sizes $[27,128]$ and [128,41412][128,41*41*2]. Mish activations are applied after each intermediate layer except the last projection layer.
  • CAMEL: The architecture consists of two networks, V-Net and C-Net
    $G(f(x);\theta,w)=c((f(x);\theta)+w^{\top}v(f(x);\theta).$ (B.1)
    The score network $V_{\phi}$ is a 5-layer MLP with width 320 and Tanh activations, with layer sizes $[d,320,320,320,320,r]$ The coefficient network $c_{\theta}$ is a 3-layer MLP with width 128 and Tanh activations, with layer sizes $[d,128,128,1]$,, where $d$ is the size of the flatten $f(x)$, and $r=5$
  • SMDP: Since we do not have an explicit physical operator to evaluate $P^{-1}(z,f(x))$, we use only the score network, implemented as a 6-layer MLP with layer sizes $128,512,256,128,64,32,d_{z}]$, where $d_{z}$ denotes the dimension of the latent space.
  • PIANO: We use the ground-truth material properties as the physical-invariant embedding. The loading field is first processed by a lifting layer, while the material parameters are passed through an attention module implemented as a 3-layer MLP with hidden dimension 32. The outputs of the lifting layer and the attention module are then concatenated and fed into a convolutional layer with width 26, followed by a 3-layer IFNO with width 26 and modes 8. For both the forward and inverse solution procedures, we adopt the same settings as in the FNO baseline.
  • InVAErt: We directly take the InVAErt implementation from Lingsch et al. (2024) and define the encoder, the VAE encoder and the decoder as 4-layer MLPs with hidden dimension 96 and silu activation function.
  • FUSE: For the forward model, we take three Fourier layers in addition to the first band-limited lifting layer that increases the dimension of the parameters and performs an inverse Fourier transform. On the other hand, the inverse model maps the functional input to the parameter space by employing a concatenation of two Fourier layers and a band-limited forward Fourier transform that generates a fixed-size latent representation of the input function, followed by a flow-matching posterior estimation (FMPE) that maps to the final parameter output. All Fourier layers have a latent width of 32 and 8 modes, while the FMPE flow has 4 layers of width 360.
  • FUSE-f: Since the original FUSE model assumes a constant loading field, it cannot handle situations where the input loading field changes that create multiple instances of a PDE system. We therefore create a FUSE variation (denoted as “FUSE-f”) that takes a concatenation of both the displacement field and the loading field as input to the inverse model. For the forward model, we concatenate the loading field to the output of the first band-limited lifting layer and subsequently use a one-layer MLP to map it back to the original dimension. All other settings are the same as the original FUSE baseline.
@@
  • convVAE: We use a convolutional layer with 136 kernels of size $3\times 3$ with a stride of 2 pixels and a fully connected layer of size [59976, 5] as the encoder, and a fully connected layer of size of [5, 59976] and a transposed convolutional layer with 2 kernels of size $3\times 3$ as the decoder.
@@
Model-agnostic architecture. The proposed DisentangO architecture is model-agnostic and can incorporate any neural operator as the forward solver, including FNO, UFNO, and WNO. This flexibility stems from DisentangO’s design, which only requires designating the lifting layer as the task-wise adaptation layer and adding the VAE structure for latent disentanglement (i.e., the encoder and the first decoder in Figure 1). DisentangO’s overall performance scales directly with the underlying neural operator’s capabilities. As long as the chosen NO achieves comparable forward prediction accuracy to IFNO, DisentangO will maintain similar performance, since (1) the upper bound of forward prediction accuracy depends on the base neural operator, and (2) the inverse prediction accuracy depends on how well the latent factors are encoded in the task-wise lifting layer parameters. We select IFNO for three key advantages: (1) parameter efficiency: the layer-independent parameter setting significantly reduces trainable parameters compared to alternatives, (2) theoretical guarantees: IFNO is a provably universal approximator as PDE solution operators, and (3) natural integration: seamless compatibility with MetaNO’s lifting layer approach for task-wise adaptation. If the IFNO part is replaced with other neural operators such as the considered baselines of FNO, UFNO, and WNO as in Table 1, the forward prediction accuracy will decrease (cf. the data test errors in Table 1), meaning that the upper bound of DisentangO’s forward prediction accuracy will decrease. Additionally, the inverse prediction accuracy (i.e., the z test error in Table 1) will likely drop as well, because the degraded forward prediction accuracy typically indicates degraded encoding of latent information in the lifting layer, which will negatively impact the inverse modeling even if the VAE part can reconstruct the lifting layer parameters perfectly. As IFNO performs the best in forward prediction and requires the least amount of trainable parameters, we choose IFNO as the forward prediction backbone in DisentangO.
To demonstrate DisentangO’s compatibility with alternative neural operators, we replace IFNO with UFNO in the DisentangO architecture. The performance comparison is presented in Table 4.
Table 4: Test errors and the number of trainable parameters in experiment 1. Bold number highlights the best method.
Models #param per-epoch Test errors
time (s) data $\bm{z}$ (SC1)
DisentangO 697k 12.2 1.65% 4.63%
MetaNO 296k 9.8 1.59% -
MetaUFNO 304k 11.4 3.55% -
DisentangO-UFNO 688k 14.2 6.61% 9.85%
UFNO 720k 21.2 7.61% 11.23%
@@ B.3 Mechanical MNIST benchmark
We demonstrate the latent traversal based on a randomly picked loading field as DisentangO-2 input. One can clearly see that the digit changes from “6” to “0” and then “2” from the top left moving down, and from “6” to “1” and then “7” moving to the right. Other digits are visible as well such as “7”, “9”, “4” and “8” in the right-most column. This corresponds well to the distribution of the latent clustering in Figure 2. We also provide two exemplary MMNIST latent interpretations of the learned DisentangO-15 in Figures 7 and 8.
@@
Figure 7: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 10 encodes the information for digit ‘8’, as is evidenced by the fact that all y<0.4y<-0.4 regions on the scatterplots contain only digit ‘8’.
@@
Figure 8: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 12 encodes the information for digit ‘3’, as is evidenced by the fact that all y<0.45y<-0.45 regions on the scatterplots contain only digit ‘3’.
B.4 Latent visualization and physical interpretation
As the hidden parameter field $b^{\eta}$ of the PDE is generally not accessible, especially in the semi-supervised or unsupervised settings, one cannot directly reconstruct $b^{\eta}$ from the learned latent variables $z$. Through disentangled representation learning, the latent variables $z$ are anticipated to contain critical information of $b^{\eta}$, but there does not exist a direct and explicit mapping from $z$ to $b^{\eta}$. As a result, one cannot directly reconstruct $b^{\eta}$ from $z$. However, there are several tricks one can play with to obtain meaningful interpretations from the learned $z$. On one hand, one can feed a randomly selected input into DisentangO and manually define a desired $z$ in the latent space and perform a forward pass using the trained DisentangO. $b^{\eta}$ can then be visualized from the output, as demonstrated in Figure 6 in the MMNIST experiment. On the other hand, one can also train a simple MLP and construct a mapping from $z$ to $b^{\eta}$, provided that $b^{\eta}$ is available. With this mapping at hand, one can then traverse $z$ and visualize what each dimension of $z$ controls. This is demonstrated in Figure 5 in the third experiment.
B.5 Further discussion
Probabilistic latent space. We use a variational (probabilistic) encoder rather than a deterministic map for two reasons:
  • Regularization and stability of the inverse map. The inverse operator in ill-posed PDE settings is highly sensitive: small perturbations in the solution $u$ can correspond to large changes in the inferred parameters. The VAE’s KL term imposes a distributional prior on the latent variables, preventing the encoder from collapsing to arbitrarily sharp or unstable mappings. This yields a smoother, more stable inverse that generalizes better under noise and limited data.
  • Identifiability and disentanglement. The identifiability results we reference (e.g., up to component-wise invertible transformations) rely on the latent distribution having a simple, factorized prior. The VAE provides exactly this structure: a normalized, independent latent prior that constrains the representation and supports the theoretical guarantees.
Scalability. Scaling interpretable neural operators to complex, high-dimensional PDEs is critical for broader impact. Our work is structured to first establish the foundational capabilities of DisentangO in controlled settings where disentanglement and representation quality can be rigorously verified. Notably, our work tackles scenarios with partially or fully unknown physics: the second experiment demonstrates DisentangO’s ability to disentangle and identify latent codes when only partial knowledge is available, and the third experiment aims to mimic a real-world fully unsupervised scenario where the physical parameters are entirely unknown, yet DisentangO successfully extracts interpretable latent factors. DisentangO is designed to be inherently scalable for several key reasons: (1) Modular architecture: DisentangO inherits the computational efficiency of the underlying neural operator backbone (IFNO in our implementation), which supports batched, GPU-accelerated inference in high dimensions. (2) Constant latent dimension: the latent code dimensionality remains low regardless of spatial resolution or system complexity, enabling practical inverse inference even for complex systems. (3) Linear scaling: the latent inference network scales linearly with the number of PDE instances, independent of spatial dimensionality. (4) Plug-and-play design: DisentangO can incorporate any state-of-the-art neural operator as its backbone, inheriting its scalability properties. In essence, the scalability of DisentangO is governed by the scalability of the neural operator used for forward modeling. As long as the NO component performs well and the training data spans sufficient variability in physical system parameters, DisentangO remains applicable.
Encoding $b$ in lifting layers. Our choice to encode the parameters $b$ in the lifting layers $\theta_{P}$ is justified through three complementary perspectives. (1) Theoretical expressivity: as demonstrated in MetaNO (Zhang et al., 2023), varying only the lifting layer across tasks is sufficient to universally approximate a wide class of parametric operators. DisentangO inherits this universal approximation property while maintaining a meta-learning framework where the core iterative Fourier layers serve as a shared meta-operator across tasks. (2) Representation identifiability: restricting adaptation to the lifting layer significantly enhances the identifiability of latent representations, a critical requirement for disentangled inverse learning. When adaptation is allowed in deeper layers, the mapping between latent codes and operators becomes more diffuse and less interpretable. (3) Empirical validation: our experiments across diverse PDE types confirm that this design choice enables DisentangO to: (a) accurately model task-specific operators, (b) recover interpretable physical factors, and (c) generalize to unseen parameter combinations and interpolation tasks.
Interpretability beyond parameter reconstruction. Our approach provides interpretability at the level of identifiable physical factors, which is often more meaningful than raw parameter reconstruction. Many physical parameters are inherently non-identifiable from observational data alone (e.g., material properties in regions with zero stress). Our method identifies the factors that actually govern system behaviors. Moreover, our experiments demonstrate clear interpretable patterns: (1) MMNIST (Figure 6): latent traversal reveals smooth transitions between digit patterns, showing the method captures meaningful morphological variations, (2) Heterogeneous materials (Figure 5): each latent dimension controls specific physical aspects: border rotation, relative fiber orientation, and segment-wise fiber orientation, (3) Semi-supervised results (Figure 2): latent factors successfully cluster different digit classes without explicit supervision. Our theoretical results on component-wise identifiability (Theorems 1 and 2) guarantee that the learned factors $\hat{z}$ correspond to true generative factors $z$ up to invertible transformations. This provides theoretical grounding for interpretability claims. In terms of practical utility, even without exact $b$ reconstruction, the disentangled factors can enable system classification and similarity assessment, design space exploration through latent traversal, transfer learning to new parameter regimes, and anomaly detection in physical systems.
Quantitative interpretability metric. The evaluation of interpretability depends critically on the availability of ground truth, which varies across our three practical scenarios, i.e., supervised, semi-supervised, and unsupervised. In the fully supervised setting corresponding to experiment 1, When true generative factors are available, we use quantitative metrics. The interpretability metric is the latent supervision test error (cf. Table 1), which directly measures how well the learned latent variables $z$ recover the true physical parameters $b$. In the semi-supervised setting corresponding to experiment 2, the goal of interpreting physics is to discover the underlying microstructure governing the deformation. Under this setting, since the true generative factors are not available, one cannot use any closed-form metric to evaluate interpretability. However, since the ground-truth microstructure generation is controlled by MNIST digits, successful disentanglement should reveal this digit information. We thus follow the standard evaluation methods in disentangled representation learning and investigate if the learned latent variables indeed contain the digit information via latent traversal, and Figure 6 confirms that our learned latent variables capture the underlying digit structure, demonstrating physical interpretability through microstructure discovery. In the third setting of unsupervised learning, without ground truth labels, we evaluate interpretability through latent traversal analysis. Figure 5 demonstrates that our method discovers meaningful physical factors: border rotation between segments, relative fiber orientation between segments, and fiber orientation of individual segments—all physically meaningful properties for understanding the microstructure mechanism.
Physical insights across different PDE types. Our method is designed for parametric PDE families—scenarios where underlying physics follows the same governing equations with varying parameters. This represents a very common and scientifically important class of problems in computational physics and engineering. While our current framework focuses on parametric variation within PDE families, the core architectural principles can extend to different PDE types. For example, in multi-physics systems, our approach could handle different but related physics (e.g., heat conduction vs. diffusion-reaction) by learning shared latent factors capturing common physical principles while using task-specific decoders. Extending to heterogeneous PDE types represents an exciting research frontier that would require modular architectures for different equation types, shared representation learning across physics domains, and physics-informed constraints for cross-domain transfer.
Accuracy-interpretability trade-off. As discussed, MetaNO represents the forward prediction upper-bound for Disentango. The accuracy gap between the two is primarily due to insufficient latent dimensions relative to data complexity. Taking experiment 3 for instance, the data generation process may involve more than 30 true generative factors due to numerical solver noise and system complexity. Table 3 clearly shows the error decreases from 25.18% to 5.28% as we increase latent dimensions from 2 to 30—a trend that would continue with higher dimensions. We stopped at 30 dimensions because: (1) 5.28% error is reasonably accurate for practical use, and (2) limited computational resource. In practice, users should gradually increase latent dimensions until convergence to MetaNO’s performance, ensuring the model capacity matches the true generative factors. DisentangO prioritizes joint modeling with interpretability over pure forward accuracy. For applications requiring only forward prediction, specialized NOs remain more accurate. However, for scientific discovery where understanding mechanisms is essential, the slightly degraded accuracy with full interpretability represents a favorable trade-off.
@@ Appendix C Data Generation
C.1 Experiment 1 - Supervised forward and inverse PDE learning
@@
where $\langle\cdot\rangle$ indicates the Macaulay bracket, $\alpha$, $k_{1}$ and $k_{2}$ are the fiber angle, modulus and exponential coefficient, respectively, $E$ denotes the Young’s modulus of the matrix, $\nu$ is Poisson’s ratio, and $S(\alpha)$ describes the fiber strain of the two fiber groups, $S(\alpha)=\frac{\bar{I}_{4}(\alpha)-1+|\bar{I}_{4}(\alpha)-1|}{2}$. $\bar{I}_{i}$ is the $i^{\text{th}}$ invariant of the right Cauchy-Green tensor $\bm{C}$, $\bar{I}_{1}=tr(\bm{C})$ and $\bar{I}_{4}=\bm{n}^{T}(\alpha)\bm{C}\bm{n}(\alpha)$, with $\bm{n}(\alpha)=[\cos(\alpha),\sin(\alpha)]^{T}$. In this context, the data generation process is controlled by sampling the material set $\{E,\nu,k_{1},k_{2},\alpha\}$, and the latent factors can be learned consequently in a supervised fashion. The physical parameters are sampled from $\frac{k_{1}}{k_{2}}\sim\mathcal{U}[90,100]$ $k_{2}\sim\mathcal{U}[0.001,0.1]$, $E\sim\mathcal{U}[0.5001,0.6001]$, $\nu\sim\mathcal{U}[0.2,0.3]$, and $\alpha\sim\mathcal{U}[\pi/10,\pi/2]$. To generate the high-fidelity (ground-truth) dataset, we sample 220 material sets, which are split into 200/10/10 for training/validation/test, respectively. For each material set, we sample $50$ different vertical traction conditions $T_{y}(\bm{x})$ on the top edge from a random field, following the algorithm in Lang & Potthoff (2011b); Yin et al. (2022). $T_{y}(\bm{x})$ is taken as the restriction of a 2D random field, $\phi(\bm{x})=\mathcal{F}^{-1}(\gamma^{1/2}\mathcal{F}(\Gamma))(\bm{x})$, on the top edge. Here, $\Gamma(\bm{x})$ is a Gaussian white noise random field on $\mathbb{R}^{2}$, $\gamma=(w_{1}^{2}+w^{2}_{2})^{-\frac{2}{4}}$ represents a correlation function, and $w_{1}$, $w_{2}$ are the wave numbers on $x$ and $y$ directions, respectively. Then, for each sampled traction loading, we solve the displacement field on the entire domain by minimizing the potential energy using the finite element method implemented in FEniCS (Alnæs et al., 2015). Sample data of the obtained dataset is illustrated in Figure 9. In this setting, the model takes as input the padded traction field and learns to predict the resulting displacement field.
@@
C.2 Experiment 2 - Mechanical MNIST benchmark
@@
C.3 Experiment 3 - Unsupervised heterogeneous material learning
@@
Varying fiber orientation distribution. We generate the samples by controlling the predefined parameters of Gaussian Random Field (GRF). With the GRF sharpened by the thresholding, the values to binary field represent two distinct fiber orientations. The binary field is smoothed using a windowed convolution. To address boundary conditions, the matrix is padded with replicated edge values, ensuring that the convolution works uniformly across the entire grid. After the smoothing process, the padded sections are removed, and the remaining field is used to construct the fiber field. We use two fixed fiber angles $\frac{\pi}{3},\frac{2\pi}{3}$ for the corresponding binary field. We generate 300 material sets, each with 500 loading/displacement pairs, and divide these into training, validation, and test sets in a 200/50/50 split. Exemplar samples from this dataset are illustrated in Figure 11. These samples demonstrate the variability of $\bm{b}$, which is critical for the latent variable identifiability in unsupervised learning settings, as proved in Theorem 2.
Varying fiber orientation magnitude and segmentation line rotation. Instead of controlling the fiber orientation angles on the binary field as two constant values, we sample the orientation distribution consisting of two segments with orientations $\alpha_{1}$ and $\alpha_{2}$ on each side, respectively, separated by a line passing through the center. The values of $\alpha_{1}$ and $\alpha_{2}$ are independently sampled from a uniform distribution over $0,2\pi]$, and the centerline’s rotation is sampled from $0,\pi]$. We generate 300 material sets, each with 500 loading/displacement pairs, and divide these into training, validation, and test sets in a 200/50/50 split.
@@
The loading in this example is taken as the body load, $\bm{f}(\bm{x})$. Each instance is generated as the restriction of a 2D random field, $\phi(\bm{x})=\mathcal{F}^{-1}(\gamma^{1/2}\mathcal{F}(\Gamma))(\bm{x})$. Here, $\Gamma(\bm{x})$ is a Gaussian white noise random field on $\mathbb{R}^{2}$, $\gamma=(w_{1}^{2}+w^{2}_{2})^{-\frac{5}{4}}$ represents a correlation function, $w_{1}$ and $w_{2}$ are the wave numbers on $x$ and $y$ directions, respectively, and $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and its inverse, respectively. This random field is anticipated to have a zero mean and covariance operator $C=(-\Delta)^{-2.5}$, with $\Delta$ being the Laplacian with periodic boundary conditions on $0,2]^{2}$, and we then further restrict it to $\Omega$. For the detailed implementation of Gaussian random field sample generation, we refer interested readers to Lang & Potthoff (2011a). Then, for each sampled loading field $\bm{f}_{i}(\bm{x})$ and microstructure field $\bm{b}^{\eta}(\bm{x})$, we solve for the displacement field $\bm{u}^{\eta}_{i}(\bm{x})$ on the entire domain.
C.4 Experiment 4 - Synthetic dataset for identifiability demonstration
We demonstrate using experiment 4 that the disentangled latent factors from DisentangO correspond to the true generative factors. We generate synthetic data following the process in equation C.2. We work with latent variables $\bm{z}$ of dimension 2 and sample from $\bm{z}\sim\mathcal{N}(\mu_{\bm{b}},\sigma_{\bm{b}}^{2}\mathbf{I})\in\mathbb{R}^{2}$ where for each microstructure $\bm{b}$, we sample $\mu_{\bm{b}}\sim\mathcal{U}[-4,4]$ and $\sigma_{\bm{b}}\sim\mathcal{U}[1,10]$ and $f\sim\mathcal{U}[0,1]\in\mathbb{R}^{3}$, $M\sim\mathcal{U}[0,1]\in\mathbb{R}^{2\times 2}$, $A=[I,2I,3I,4I]^{T}$, $M\in\mathbb{R}^{8\times 2}$, $b\sim\mathcal{U}[0,1]\in\mathbb{R}^{8}$, $W\sim\mathcal{U}[0,1]\in\mathbb{R}^{4\times 1}$, $l\sim\mathcal{U}[0,1]\in\mathbb{R}$,
[…]
We generate 700 data pairs of $(z,\theta)$ corresponding to 700 tasks, and for each task, we generate 200 samples of $(f,u)$ pairs.
$d$ 3 5 7 9
MCC 0.7836 0.8013 0.8237 0.9121
$R^{2}$ 0.5673 0.6026 0.6473 0.8242
Rank(w)(w) 3 4 4 4
Table 5: The MCC and $R^{2}$ scores to evaluate identifiability on the synthetic dataset in experiment 4.
Figure 12: The scatter plot of the true generative variables $z$ and the disentangled latent variables $\hat{z}$ from the synthetic dataset with $d=9$ in experiment 4.
We conduct experiments on the synthetic dataset and split the tasks into dd/100/500 for training, validation and test, respectively, with $d$ chosen from the set of $(3,5,7,9)$ to investigate the effect of the number of training tasks on the identifiability. In particular, we measure the component-wise identifiability of the latent variables $\bm{z}$ by computing the Mean Correlation Coefficient (MCC) and the coefficient of determination $R^{2}$ scores on the test dataset, and report the results in Table 5, where we observe a monotonic growth on both MCC and $R^{2}$ with the increase of the number of training tasks $d$. Figure 12 illustrates the alignment on the test dataset between the true generative factors $z$ and the disentangled factors $\hat{z}$ from DisentangO in the case of $d=9$. To verify Assumption 4, we calculate the rank of the matrix $w(\bm{z},\bm{b})$. Specifically, for each $\theta$, with the VAE encoder, we calculate the $\mu_{\bm{b}}$ and $\sigma_{\bm{b}}$, sample $z$, and calculate the corresponding $\dfrac{\partial q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}}$ and $\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}^{2}}$, where $i=1,\cdots d_{z},j=1\cdots d$.